OpenStudy (anonymous):

how do i solve x^2+8x+15 divided by x+5

OpenStudy (anonymous):

u can't solve something that doesn't equal to something else

OpenStudy (anonymous):

u just posted an expression.

OpenStudy (anonymous):

Its long division with polynomials

OpenStudy (anonymous):

factor and cancel

OpenStudy (anonymous):

\[\frac{x^2+8x+15}{x+5}=\frac{(x+5)(x+3)}{x+5}=x+3\]

OpenStudy (anonymous):

Im not understanding. I have to divide . which doesn't require factoring.

OpenStudy (anonymous):

you mean like long divistion?

OpenStudy (anonymous):

____________ x+5 { x^2 + 8x + 15

OpenStudy (anonymous):

divide. x goes in to x^2 x times, multuply, subtract, etc just like with whole numbers

OpenStudy (anonymous):

yes